OpenStudy (anonymous):

WILL FAN AND MEDAL 1+cos(12x)=? A.2cos^2(12x) B.2sin^2(12x) C.2cos^2(6x) D.2sin^2(6x)

OpenStudy (anonymous):

Here's a really good list of trig identities http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (xapproachesinfinity):

hmm we have \(\cos(12x)=\cos(2(6x))=\cos^26x-\sin^26x\) then \(1+\cos(12x)=1+\cos^26x-\sin^26x=2\cos^26x\)

OpenStudy (xapproachesinfinity):

the identity i used is \(\cos2x=\cos^2x-\sin^2x\)

OpenStudy (jhannybean):

Awesome job @xapproachesinfinity

OpenStudy (anonymous):

Another way is to see that \[\Large \frac{ 1+\cos(2x) }{ 2 }=\cos^2x\]

OpenStudy (anonymous):

You're only left with C because \[\Large \frac{1-\cos(2x)}{2}=\sin^2x\] and there is no subtraction sign. Anyway, just to be sure set C equal to 1+cos(12x) \[\Large 1+\cos(12x)=2\cos^2(6x)\] Divide both sides by two \[\Large \frac{1+\cos(12x)}{2}=\cos^2(6x)\] And you're left with your identitity